Operations Report – May 16th

Crew 194 Operations Report 16.05.18

SOL: 3

Name of person filing report: Ben Easter

Non-nominal systems: Power per prior documentation

Notes on non-nominal systems: Continued issues as already described

Generator (hours run): Overnight approximately 10

Solar— SOC 0%

Diesel – Over 50%

Propane – psi. Not checked

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 75%

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 134955

Toilet tank emptied: No
Date last emptied: By director prior to our arrival

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 3 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs:

Deimos rover used: No
Hours: N/A
Beginning charge: 100%
Ending charge: XX %
Currently charging: Yes or No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Hours: 35.8
Beginning charge: 100 %
Ending charge: 90%
Currently charging: Yes

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 94 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: XX %
Currently charging: Yes or No

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: working well

Summary of suits and radios: working well

Summary of Hab operations: Intermittent gas odor with turning on water heater.

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Daily summary – May 16th

Crew 194 Daily Summary Report 16.05.2018

MDRS Daily Summary Report for sol 3

Summary Title

A Successful Rocket Launch!

Mission Status:


Sol Activity Summary:

The majority of our day was spent preparing for successful launch of a high powered rocket. We obtained final clearance from the FAA earlier this week, and called in our NOTAM to clear the airspace. Last night and this morning, we performed final preparations on the rocket and then assembled the launch tower during EVA. We had a successful launch, which was incredible to see! We are planning to recover the rocket this evening after it cools down, and will see what the pictures look like.

Look Ahead Plan: We will continue our learning with discussion of ultrasound in spaceflight as well as genitourinary issues.

Anomalies in work:


Weather:Sunny, Very Warm

Crew Physical Status:Very happy after the launch this morning!

EVA: A successful launch!

Reports to be file:

Operations Report

Journalist Report

EVA Report

Support Requested:


Copyright © The Mars Society. All rights reserved. | Main Site